One year ago, you invested $2,700. Today, it is worth $3,500. What rate of interest did you earn?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

3500=2700*(1+r/100)

(3500/2700)=(1+r/100)

(1+r/100)=1.2963

r=1.2963-1

=29.63%(Approx).